Evaluate the following indefinite integrals. Be sure to clearly identify your u and du.
5. ſcot (90) do
Answer #1

5)

\small \int cot(9x)dx

Substitute \small u=9x\rightarrow du=9dx

\small \Rightarrow dx=\frac{du}{9}

The integration becomes  \small \int cot(u)\frac{du}{9}

I cot (uldu 9

We need to find

\small \int cot(u)du

\small =\int \frac{cos(u)}{sin(u)}du

Substitute \small v=sin(u)\rightarrow dv=cos(u)du

So, integration becomes

\small \int \frac{1}{v}dv=ln|v|+c

\small =ln|sin(u)|+c   \small (\because v=sin(u))

Putting the value in the original integral, we get

\small \frac{1}{9}\int cot(u)du=\frac{1}{9}ln|sin(u)|+c

\small =\frac{1}{9}ln|sin(9x)|+c   \small (\because u=9x)

\small \therefore \int cot(9x)dx=\frac{1}{9}ln|sin(9x)|+C, where C is a constant.